OpenStudy (anonymous):

Use the Law of Levers to solve the problem. Three friends pay $100 each to share a cottage on a camping trip. The cost per person varies inversely as the number of people on the camping trip. How many people would have to be on the trip to make the cost $60 per person? Answer 6 people 4 people 5 people How would you set up this problem?

OpenStudy (perl):

law of levers?

OpenStudy (anonymous):

physics http://physics.weber.edu/carroll/archimedes/lever.htm

OpenStudy (nurali):

100/60 = x/3. 100*3=60x 300=60x x=300/60 x=5

OpenStudy (anonymous):

Thank you! That's how I set it up too I just wanted to make sure I wasn't completely doing it wrong.

OpenStudy (nurali):

My Pleasure.

OpenStudy (perl):

how does that use levers?

OpenStudy (perl):

@Nurali how does that use levers?

OpenStudy (perl):

what the flutter does this have to do with levers